For a class demonstration, your physics instructor pours 1.00 kg of steam at 100
Answer #1

Let the final temperature be T.

heat lost by steam = 1 * 2.26 * 10^6 + 1* 4186 * (100-T)

heat gaiend by ice = 5 * 3.33 * 10^5 + 5 * 4186 * T

These are equal which gives:

22.6 * 10^5 - 16.65 * 10^5 - 418600 = 6 * 4186 * T

so T = 7.023 degrees
Answer #2

Here is what I solved before, please modify the figures as per your question. Please let me know if you have further questions. Ifthis helps then kindly rate 5-stars.

For a class demonstration, your physics instructor pours 1.25kg of steam at 100.0
Answer #3

heat gained by ice = heat lost by steam

m_ice*Lf + m_ice*C_water*(T-0) = m_steam*Lv + m_steam*C_water*(100-T)


T*(m_ice*C_water + m_steam*C_water) = -m_ice*Lf + m_steam*Lv + m_steam*C_water*100

T = (-m_ice*Lf + m_steam*Lv + m_steam*C_water*100)/(m_ice*C_water + m_steam*C_water)

= (-5*3.33*10^5 + 1*2.26*10^6 + 1*4186*100)/(5*4186 + 1*4186*100)

= 2.306 degrees celsius